Here are measurements from my sample,
L1 (lowest) hotspot :
CCT = 4114K (Duv 0.0027) CRI (Ra) = 94.5 [ R9 = 79.4 ]
IES TM-30-15 Rf = 91.30 Rg = 98.94
L2 :
CCT = 4138K (Duv 0.0028) CRI (Ra) = 94.1 [ R9 = 78.3 ]
IES TM-30-15 Rf = 91.44 Rg = 99.25
L3 :
CCT = 4147K (Duv 0.0025) CRI (Ra) = 93.5 [ R9 = 76.5 ]
IES TM-30-15 Rf = 91.57 Rg = 99.72
L4 :
CCT = 4192K (Duv 0.0014) CRI (Ra) = 93.0 [ R9 = 75.2 ]
IES TM-30-15 Rf = 91.30 Rg = 100.53
L5 (max) :
CCT = 4221K (Duv –0.0005) CRI (Ra) = 92.8 [ R9 = 74.0 ]
IES TM-30-15 Rf = 90.74 Rg = 101.15
L5 ~corona : CCT = 4037K (Duv 0.0011)
L5 ~spill : CCT = 4447K (Duv –0.0045)
TM-30-18 report :

At first I was surprised to see low Duv numbers from my measurements, and then noticed that the bezel wasn’t fully screwed from previous tinkering, here are the measurement with the bezel slightly unscrewed (~0.8mm gap) :
L5 hotspot (unscrewed) : CCT = 4270K (Duv –0.0039)
I imagine the TIR must catch less greener light from the sides of the dome.
The bare LED itself is quite rosy directly above it :
L5 bare led (directly above) : CCT = 4375K (Duv –0.0076)